ご14.jass) In Mostar, Bosnia, the ultimate test of a young mans courage once was to jump off a 400- 4. ,45s) In Mostar, Bosnia, the ultimate test of a young mans courage once was to jump off a 400- yeár-old bridge (now destroyed) into the River Neretva, 23.0 m below the bridge. (a) How long did the jump last? (b) How fast was the diver traveling upon impact with the water? (c) If the speed of sound in air is 340 m/s, how long after the diver took off did a spectator on the bridge hear the splash?
Answer #1

14)h = 23 m

We know that

s = ut + 1/2 at^2

here, s = h = 23 m ; a = g = 9.8 m/s^2 ; t needs to be calculated

h = 1/2 g t^2

t = sqrt (2h/g)

t = sqrt (2 x 23/9.8) = 2.17 s

Hence, t = 2.17 s

b)from conservation of energy

KE = PE

1/2 m v^2 = m g h

v = sqrt (2 g h)

v = sqrt (2 x 9.8 x 23) = 21.23 m/s

Hence, v = 21.23 m/s

c)We know that

v = d/t => t = d/v

t = 23/340 = 0.068 s

Hence, t = 0.068 s